OpenStudy (anonymous):

Find x and y in simplest radical form. x = 5√3; y = 10 x = 5√3; y = 5 x = 10√3; y = 5 x = 10√3; y = 10

OpenStudy (anonymous):



OpenStudy (anonymous):

@Auxuris @mikurout

OpenStudy (anonymous):

@eric_d

OpenStudy (mikurout):

You can use sin(of an angel ) or cosine formula.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

im not very good at math

OpenStudy (mikurout):

Ok sorry, use tan 30 = p/b

OpenStudy (anonymous):

hmmm....ugh im lost Im soooo sorry i just dont know what im supposed to do next

OpenStudy (mikurout):

tan 30 = \[\sqrt{3}\]/ 3

OpenStudy (anonymous):

so does that equal 10?

OpenStudy (mikurout):

\[\frac{ 5 }{x ? }\] = \[\frac{ \sqrt{3} }{ ?3 }\]

OpenStudy (mikurout):

so x = \[5\sqrt{3}\]

OpenStudy (anonymous):

how do you solve that?

OpenStudy (mikurout):

do you know this thing? tan@ = p/b ?

OpenStudy (anonymous):

no :(

OpenStudy (mikurout):

any idea about trigonometric functions?

OpenStudy (anonymous):

no, i dont think ive learned that yet

OpenStudy (mikurout):

Have you learned about pythagoras theorem?

OpenStudy (anonymous):

yes

OpenStudy (mikurout):

without using trigonometric function you can't solve this question. but using pythagoras theorem you can chose the right option from the given four options.

OpenStudy (mikurout):

just use that theorem and put the values of x and y . After it you will get the right ans.

OpenStudy (anonymous):

im being totally honest, i dont know what to do, can you walk me through

OpenStudy (mikurout):

sure

OpenStudy (anonymous):

thank you! ;)

OpenStudy (nincompoop):

I don't understand the drawing is that a 90 degree angle?

OpenStudy (anonymous):

yes

OpenStudy (mikurout):

according to your figure (square of) 5 + (square of) x = (square of ) y . is that clear?

OpenStudy (nincompoop):

do you need to learn the vocabulary first before starting out with any solution? I think it will serve you a great deal in the future that is when to use sine, cosine and tangent

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

many people jump into solutions without considering what you already know or do not yet know.

OpenStudy (anonymous):

ok ill keep that in mind :)

OpenStudy (nincompoop):

let us follow a certain theme here okay? let us consider the "angle of reference," that is which angle are we looking at and other parts associated with it.

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):